A householder is worried about the radiation emitted by the granite worktop in his kitchen.

1 kg of granite has an activity of 1250 Bq. The kitchen worktop has a mass of 180 kg.

Calculate the activity of the kitchen worktop in Bq.




1
Expert's answer
2021-02-17T10:41:22-0500

The activity is an additive quantity. If 1 kg of granite has an activity of 1250 Bq, then 180 kg will have the following activity:


"A = 180kg\times1250\dfrac{Bq}{kg} = 2.25\times 10^5Bq"

Answer. "2.25\times 10^5Bq".